A) The following data lists the average monthly snowfall for January in 15 cities around the US. I need to find the mean, variance, and standard deviation.
36 12 6 18 7 36 31 4 12 23 27 14 3 25 47
B)A math test has a mean of 49 and standard deviation of 6.0. I need to the the corresponding z scores for: a test score of 36 and a test score of 78.

Answer by ewatrrr(10682)


Hi
A) entering data in an Excel column and using the data analysis tool
mean = 20.0667 , varience = 180.2095, standard deviation = 13.4242
B) z = (36-49) /6 and z = (78-49)/ 6